# On the Expressive Power of a Variant of the Looped Transformer

### v1

@@ 3.2 AR(q) with Representation
We consider the autoregressive model with representation. The dynamical (time series) system is generated by
$\bm{x}_{t+1}=\bm{A}\Phi^{*}\left([\bm{x}_{t+1-q},\cdots,\bm{x}_{t}]\right)+\bm{\epsilon}_{t},$
where $\Phi^{*}\left(\cdot\right)$ is a fixed representation function (e.g., we take the L-layer MLPs), and the weight $\bm{A}$ vary from different prompts. In standard AR(q) (multivariate autoregressive) models, the representation function $\Phi^{*}\left(\cdot\right)$ is identity. Here, we investigate a more challenging situation in which the representation function is fixed but unknown. A well-behaved solver should first find the representation function and then translate it into a modified autoregressive model. With standard Gaussian priors on the white noise $\bm{\epsilon}_{t}$, the Bayesian estimator of the AR(q) model parameters admits $\arg\max_{\bm{A}}\prod_{t=1}^{N}f(\bm{x}_{t}|\bm{x}_{t-1},\cdots,\bm{x}_{t-q})=\arg\min_{\bm{A}}\sum_{t=1}^{N}\left\|\bm{x}_{t}-\bm{A}\Phi^{*}\left([\bm{x}_{t-q},\cdots,\bm{x}_{t-1}]\right)\right\|_{2}^{2}$, where $f(\bm{x}_{t}|\bm{x}_{t-1},\cdots,\bm{x}_{t-q})$ is the conditional density function of $\bm{x}_{t}$, given previous $q$ observations. A practical solver initially identifies the representation function and transforms the input time series into its representation, denoted as $\Phi^{*}(\bm{x}_{t})$. Then the problem is reduced to an autoregressive form. Similar to the previous subsection, we prove by construction that there exists a AlgoFormer, akin to human-designed algorithms, capable of effectively solving the given task.
@@
[…]
Remarks. The detailed proof can be found in Appendix A.2. The technical details are similar to Theorem 3.1. Additionally, the pre-transformer copies the feature from the previous $q$ tokens, utilizing $q$ heads for parallel processing.
@@ 3.3 Chain-of-Thought with MLPs
Chain-of-Thought (CoT) demonstrates exceptional performances in mathematical reasoning and text generation (Wei et al., 2022). The success of CoT has been theoretically explored, shedding light on its effectiveness in toy cases (Li et al., 2023) and on its computational complexity (Feng et al., 2023). In this subsection, we revisit the intriguing toy examples of CoT generated by leaky ReLU MLPs, denoted as CoT with MLPs, as discussed in Li et al. (2023). We begin by constructing an L-layer MLP with leaky ReLU activation. For an initial data point $\bm{x}\sim\mathcal{P}_{x}$, the CoT point $\bm{s}^{\ell}$ represents the output of the $\ell$-th layer of the MLP. Consequently, the CoT sequence $\{\bm{x},\bm{s}^{1},\cdots,\bm{s}^{L}\}$ is exactly generated as the output of each (hidden) layer of the MLP. The target of CoT with MLPs problem is to find the next state $\hat{\bm{s}}^{\ell+1}$ based on the CoT samples $\{\bm{x}_{1},\bm{s}_{1}^{1},\cdots,\bm{s}_{1}^{L},\bm{x}_{2},\cdots,\bm{x}_{N},\bm{s}_{N}^{1},\cdots,\bm{s}_{N}^{L},\bm{x}_{\text{test}},\hat{\bm{s}}^{1},\cdots,\hat{\bm{s}}^{\ell}\}$, where $\{\hat{\bm{s}}^{1},\cdots,\hat{\bm{s}}^{\ell}\}$ denotes the CoT prompting of $\bm{x}_{\text{test}}$. We establish by construction in Theorem 3.3 that the AlgoFormer adeptly solves the CoT with MLPs problem, exhibiting a capability akin to human-designed algorithms.
@@
[…]
@@
4 Discussion
In this section, we provide complementary insights to the results discussed in Section 3. Firstly, as discussed in the remark following Theorem 3.1, we construct the looped transformer that employs gradient descent to solve (regularized) multi-variate regression problems. However, in practical scenarios, the adoption of more efficient optimization algorithms is often preferred. Investigating the expressive power of transformers beyond gradient descent is both intriguing and appealing. As stated in Theorem 4.1, we demonstrate that the AlgoFormer can proficiently implement Newton’s method for solving linear regression problems. Secondly, the definition in Equation (1) implies the encoder-based transformer. In practical applications, a decoder-based transformer with causal attention, as seen in models like GPT-2 (Radford et al., 2019), may also be favored. For completeness, it is also compelling to examine the behavior of decoder-based transformers in algorithmic learning. Our findings, presented in Theorem 4.2, reveal that the decoder-based transformer block can also implement gradient descent in linear regression problems. The primary distinction lies in the fact that the decoder-based transformer utilizes previously observed data to evaluate the gradient, while the encoder-based transformer calculates the gradient based on the full data samples.
@@ 4.1 Beyond the Gradient Descent
In the designed transformer block, the pre-transformer initially processes the input token, transforming it into an optimization problem. Subsequently, the looped transformer implements iterative algorithms, such as gradient descent, to solve the optimization problem. Finally, the post-transformer aggregates and outputs the desired results. While our construction currently adopts gradient descent as the solver for optimization problems, practical scenarios often witness the superiority of certain optimization algorithms over gradient descent. For example, Newton’s (second-order) methods enjoy superlinear convergence under some mild conditions, outperforming gradient descent with linear convergence. This raises a natural question:
@@
In this section, we address this question by demonstrating that the designed AlgoFormer can also realize Newton’s method in regression problems.
Consider the linear regression problem given by:
@@
$\begin{array}[]{ll}\bm{M}_{0}=\alpha\bm{S},\text{ where }\alpha\in\left(0,\frac{2}{\left\|\bm{S}\bm{S}^{\top}\right\|_{2}}\right],\\ \vspace{0.5em}\bm{M}_{k+1}=2\bm{M}_{k}-\bm{M}_{k}\bm{S}\bm{M}_{k},\\ \vspace{0.5em}\bm{w}_{0}^{\text{Newton }}=\bm{M}_{0}\bm{X}^{\top}\bm{y},\\ \bm{w}_{k+1}^{\text{Newton }}=\bm{M}_{k+1}\bm{X}^{\top}\bm{y}.\end{array}$ (4)
As described in Söderström & Stewart (1974); Pan & Schreiber (1991), the above update scheme (Newton’s method) enjoys superlinear convergence, in contrast to the linear convergence of gradient descent. The following theorem states that Newton’s method in Equation (4) can be realized by the AlgoFormer.
@@
[…]
Remarks. The proof can be found in Appendix A.4. The pre-transformer performs preparative tasks, such as copying from neighboring tokens. The looped-transformer is responsible for updating and calculating $\bm{M}_{k}\bm{x}_{i}$ for each token $\bm{x}_{i}$ at every step $k$. The post-transformer compute the final estimated weight $\bm{w}_{T}^{\text{Newton }}$ and outputs the desired the results $\bm{w}_{T}^{\text{Newton}\top}\bm{x}_{\text{test}}$, where $T$ is the iteration number in Equations (2) and (4). In a related study by Fu et al. (2023), similar topics are explored, indicating that transformers exactly perform higher-order optimization algorithms. However, our transformer architectures differ, and technical details are distinct. Empirically, as shown in Section 5.3 and Appendix B.3, we find that the transformer is smarter compared to some human-designed algorithms, such as Newton’s method.
@@ 4.2 Decoder-based Transformer
In the preceding analysis, the encoder-based AlgoFormer (with full attention) demonstrates its capability to solve problems by performing algorithms. Previous studies (Giannou et al., 2023; Bai et al., 2023; Zhang et al., 2023a; Huang et al., 2023; Ahn et al., 2023) also focus on the encoder-based models. We opted for an encoder-based transformer because full-batch data is available for estimating gradient and Hessian information. However, in practical applications, decoder-based models, like GPT-2, are sometimes more prevalent. In this subsection, we delve into the performance of the decoder-based model when executing iterative optimization algorithms, such as gradient descent, to solve regression problems. The results are intriguing, as the decoder-based transformer evaluates the gradient based on previously viewed data from earlier tokens.
We consider the linear regression problem in Equation (3). Due to the limitations of the decoder-based transformer, which can only access previous tokens, implementing iterative algorithms based on the entire batch data is not feasible. However, it is important to note that the current token in a decoder-based transformer can access data from all previous tokens. To predict the label $y_{i}$ based on the input prompt $\bm{P}^{i}=\left[\bm{x}_{1},y_{1},\cdots,\bm{x}_{i}\right]$, the empirical loss for the linear weight at $\bm{x}_{i}$ is given by
@@
[…]
@@ 5 Experiment
In this section, we conduct a comprehensive empirical evaluation of the AlgoFormer’s performance in tackling challenging tasks, specifically addressing regression with representation, AR(q) with representation, and CoT with MLPs, as outlined in Section 2. Due to space limits, additional content has been provided in Appendix B. Before reporting the detailed experimental results, we provide an elaborate description of the experimental settings.
@@
Experimental settings. In all experiments, we adopt the decoder-based AlgoFormer, standard transformer (GPT-2), and vanilla looped transformer (Yang et al., 2024). We utilize $N=40$ in-context samples as input prompts and $d=20$ dimensional vectors with $D=256$ dimensional positional embeddings for all experiments. To ensure fairness in comparisons, all models are trained using the Adam optimizer, with 500K iterations to ensure convergence. Both the pre- and post-transformers are implemented as one-layer transformers. For task-specific settings and additional details, please refer to Appendix B.1.
Training strategy. Our training strategy builds upon the methodology introduced in Yang et al. (2024). Let $\bm{P}^{i}=[\bm{x}_{1},f(\bm{x}_{1}),\cdots,\bm{x}_{i-1},f(\bm{x}_{i-1}),\bm{x}_{i}]$ represents the input prompt for $1\leq i\leq N$. We denote the AlgoFormer as $\text{TF}_{\text{Algo}}^{t}(\cdot;\bm{\Theta})$, where $f(\cdot)$ is a task-specific function and $t$ indicates the number of loops (iterations) in Equation (2), and $\bm{\Theta}$ represent the transformer parameters. Instead of evaluating the loss solely on $\text{TF}_{\text{Algo}}^{T}(\cdot;\bm{\Theta})$ with $T$ iterations, we minimize the expected loss over averaged iteration numbers:
where $T_{0}=\max\{T-\Delta T,0\}$. Here, the prompt formulation and the above loss may slightly differ for different tasks. For example, in the AR(q) task, the prompt is reformulated as $\bm{P}^{i}=[\bm{x}_{1},\cdots,\bm{x}_{i-1},\bm{x}_{i}]$ and $\bm{x}_{i+1}=f([\bm{x}_{i+1-q},\cdots,\bm{x}_{i}])$ is the target for prediction. But the training strategy can be easily transmitted to other tasks. Here, both the iteration numbers $T_{0}$ and $T$ are hyperparameters, which will be analyzed in Appendix B.2.
5.1 AlgoFormer Exhibits Higher Expressiveness

### v2

@@ 3.2 AR(q) with Representation
We consider the autoregressive model with representation. The dynamical (time series) system is generated by $\bm{x}_{t+1}=\bm{A}\Phi^{*}\left([\bm{x}_{t+1-q},\cdots,\bm{x}_{t}]\right)+\bm{\epsilon}_{t},$ where $\Phi^{*}\left(\cdot\right)$ is a fixed representation function (e.g., we take the L-layer MLPs), and the weight $\bm{A}$ varies from different sequences but remains constant within a single sequence. Our goal is to learn $\Phi^{*}\left(\cdot\right)$ during training and to perform in-context learning of the weight matrix $\bm{A}$. In standard AR(q) (multivariate autoregressive) models, the representation function $\Phi^{*}\left(\cdot\right)$ is identity. Here, we investigate a more challenging situation in which the representation function is fixed but unknown. A well-behaved solver should first find the representation function and then translate it into a modified autoregressive model. With standard Gaussian priors on the white noise $\bm{\epsilon}_{t}$, the Bayesian estimator of the AR(q) model parameters admits $\arg\max_{\bm{A}}\prod_{t=1}^{N}f(\bm{x}_{t}|\bm{x}_{t-1},\cdots,\bm{x}_{t-q})=\arg\min_{\bm{A}}\sum_{t=1}^{N}\left\|\bm{x}_{t}-\bm{A}\Phi^{*}\left([\bm{x}_{t-q},\cdots,\bm{x}_{t-1}]\right)\right\|_{2}^{2}$, where $f(\bm{x}_{t}|\bm{x}_{t-1},\cdots,\bm{x}_{t-q})$ is the conditional density function of $\bm{x}_{t}$, given previous $q$ observations. A practical solver initially identifies the representation function and transforms the input time series into its representation, denoted as $\Phi^{*}(\bm{x}_{t})$. Then the problem is reduced to an autoregressive form. Similar to the previous subsection, we prove by construction that there exists a AlgoFormer, akin to human-designed algorithms, capable of effectively solving the given task.
@@
There exists a designed AlgoFormer with $\text{TF}_{\text{pre}}$ (a one-layer $q$-head transformer with an $(L+1)$-layer one-head transformer), $\text{TF}_{\text{loop}}$ (a one-layer two-head transformer), and $\text{TF}_{\text{post}}$ (a one-layer one-head transformer), that predicts $\bm{x}_{N+1}$ from the data sequence $\{\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{N}\}$ by copying, transformation of the representation function and applying gradient descent for multi-variate regression. The emulation of each step is not exact, as there is some error introduced in each step. However, the error can be made arbitrarily close to zero by increasing the temperature of the softmax and adjusting another free parameter, neither of which affects the size of the network.
Remarks. The detailed proof can be found in Appendix A.2. The technical details are similar to Theorem 3.1. The input prompt to the transformer is formulated as $\bm{P}=\left[\begin{array}[]{ccccc}\bm{x}_{1}&\bm{x}_{2}&\cdots&\bm{x}_{N}\\ \bm{p}_{1}^{x}&\bm{p}_{2}^{x}&\cdots&\bm{p}_{N}^{x}\end{array}\right],$ where $\bm{p}_{i}^{x}$ denote positional embeddings and will be specified in the proof. Additionally, the pre-transformer copies the feature from the previous $q$ tokens, utilizing $q$ heads for parallel processing.
@@ 3.3 Chain-of-Thought with MLPs
Chain-of-Thought (CoT) demonstrates exceptional performances in mathematical reasoning and text generation (Wei et al., 2022). The success of CoT has been theoretically explored, shedding light on its effectiveness in toy cases (Li et al., 2023) and on its computational complexity (Feng et al., 2023). In this subsection, we revisit the intriguing toy examples of CoT generated by leaky ReLU MLPs, denoted as CoT with MLPs, as discussed in Li et al. (2023). We begin by constructing an L-layer MLP with leaky ReLU activation. For an initial data point $\bm{x}\sim\mathcal{P}_{x}$, the CoT point $\bm{s}^{\ell}$ represents the output of the $\ell$-th layer of the MLP. Consequently, the CoT sequence $\{\bm{x},\bm{s}^{1},\cdots,\bm{s}^{L}\}$ is exactly generated as the output of each (hidden) layer of the MLP. The implicit $L$-layer MLP remains the same within a single sequence but varies across different sequences, and it is learned in-context. The target of CoT with MLPs problem is to find the next state $\hat{\bm{s}}^{\ell+1}$ based on the CoT samples $\{\bm{x}_{1},\bm{s}_{1}^{1},\cdots,\bm{s}_{1}^{L},\bm{x}_{2},\cdots,\bm{x}_{N},\bm{s}_{N}^{1},\cdots,\bm{s}_{N}^{L},\bm{x}_{\text{test}},\hat{\bm{s}}^{1},\cdots,\hat{\bm{s}}^{\ell}\}$, where $\{\hat{\bm{s}}^{1},\cdots,\hat{\bm{s}}^{\ell}\}$ denotes the CoT prompting of $\bm{x}_{\text{test}}$. We establish by construction in Theorem 3.3 that the AlgoFormer adeptly solves the CoT with MLPs problem, exhibiting a capability akin to human-designed algorithms.
@@
There exists a designed AlgoFormer with $\text{TF}_{\text{pre}}$ (a seven-layer two-head transformer), $\text{TF}_{\text{loop}}$ (a one-layer two-head transformer), and $\text{TF}_{\text{post}}$ (a one-layer one-head transformer), that finds $\hat{\bm{s}}^{\ell+1}$ from samples $\{\bm{x}_{1},\bm{s}_{1}^{1},\cdots,\bm{s}_{1}^{L},\bm{x}_{2},\cdots,\bm{x}_{N},\bm{s}_{N}^{1},\cdots,\bm{s}_{N}^{L},\bm{x}_{\text{test}},\hat{\bm{s}}^{1},\cdots,\hat{\bm{s}}^{\ell}\}$ by filtering and applying gradient descent for multi-variate regression. The emulation of each step is not exact, as there is some error introduced in each step. However, the error can be made arbitrarily close to zero by increasing the temperature of the softmax and adjusting another free parameter, neither of which affects the size of the network.
@@
4 Extensions and Further Analysis
In this section, we provide complementary insights to the results discussed in Section 3. Firstly, as discussed in the remark following Theorem 3.1, we construct the looped transformer that employs gradient descent to solve (regularized) multi-variate regression problems. However, in practical scenarios, the adoption of more efficient optimization algorithms is often preferred. Investigating the expressive power of transformers beyond gradient descent is both intriguing and appealing. As stated in Theorem 4.1, we demonstrate that the AlgoFormer can proficiently implement Newton’s method for solving linear regression problems. Secondly, the definition in Equation 1 implies the encoder-based transformer. In practical applications, a decoder-based transformer with causal attention, as seen in models like GPT-2 (Radford et al., 2019), may also be favored. For completeness, it is also compelling to examine the behavior of decoder-based transformers in algorithmic learning. Our findings, presented in Theorem 4.2, reveal that the decoder-based AlgoFormer can also implement gradient descent in linear regression problems. The primary distinction lies in the fact that the decoder-based transformer utilizes previously observed data to evaluate the gradient, while the encoder-based transformer calculates the gradient based on the full data samples.
@@ 4.1 Beyond the Gradient Descent
Newton’s (second-order) methods enjoy superlinear convergence under some mild conditions, outperforming gradient descent with linear convergence. This raises a natural question:
@@
In this section, we address this question by demonstrating that the designed AlgoFormer can also realize Newton’s method in regression problems. Consider the linear regression problem given by:
@@
$\bm{M}_{0}=\alpha\bm{S},~{}\text{where}~{}\alpha\in\left(0,\frac{2}{\left\|\bm{S}\bm{S}^{\top}\right\|_{2}}\right],\hskip 5.0pt\bm{M}_{k+1}=2\bm{M}_{k}-\bm{M}_{k}\bm{S}\bm{M}_{k},\hskip 5.0pt\bm{w}_{k}^{\text{Newton}}=\bm{M}_{k}\bm{X}^{\top}\bm{y}.$ (5)
As described in Söderström & Stewart (1974); Pan & Schreiber (1991), the above update scheme (Newton’s method) enjoys superlinear convergence, in contrast to the linear convergence of gradient descent. The following theorem states that Newton’s method in Equation 5 can be realized by the AlgoFormer.
@@
There exists a designed AlgoFormer with $\text{TF}_{\text{pre}}$ (a one-layer two-head transformer), $\text{TF}_{\text{loop}}$ (a one-layer two-head transformer), and $\text{TF}_{\text{post}}$ (a two-layer two-head transformer), that implements Newton’s method described by Equation 5 in solving regression problems. The emulation of each step is not exact, as there is some error introduced in each step. However, the error can be made arbitrarily close to zero by increasing the temperature of the softmax and adjusting another free parameter, neither of which affects the size of the network.
Remarks. The proof can be found in Appendix A.4. The pre-transformer performs preparative tasks, such as copying from neighboring tokens. The looped-transformer is responsible for updating and calculating $\bm{M}_{k}\bm{x}_{i}$ for each token $\bm{x}_{i}$ at every step $k$. The post-transformer compute the final estimated weight $\bm{w}_{T}^{\text{Newton }}$ and outputs the desired the results $\bm{w}_{T}^{\text{Newton}\top}\bm{x}_{\text{test}}$, where $T$ is the iteration number in Equation 2 and Equation 5. In a related study by Fu et al. (2023), similar topics are explored, indicating that transformers exactly perform higher-order optimization algorithms. However, our transformer architectures differ, and technical details are distinct.
@@ 4.2 Decoder-based Transformer
In the preceding analysis, the encoder-based AlgoFormer (with full attention) demonstrates its capability to solve problems by performing algorithms. Previous studies (Giannou et al., 2023; Bai et al., 2023; Zhang et al., 2023a; Huang et al., 2023; Ahn et al., 2023) also focus on the encoder-based models. We opted for an encoder-based transformer because full-batch data is available for estimating gradient and Hessian information. However, in practical applications, decoder-based models, like GPT-2, are sometimes more prevalent. In this subsection, we delve into the performance of the decoder-based model when executing iterative optimization algorithms, such as gradient descent, to solve regression problems.
We consider the linear regression problem in Equation 4. Due to the limitations of the decoder-based transformer, which can only access previous tokens, implementing iterative algorithms based on the entire batch data is not feasible. However, it is important to note that the current token in a decoder-based transformer can access data from all previous tokens. To predict the label $y_{i}$ based on the input prompt $\bm{P}^{i}=\left[\bm{x}_{1},y_{1},\cdots,\bm{x}_{i}\right]$, the empirical loss for the linear weight at $\bm{x}_{i}$ is given by
@@
There exists a designed AlgoFormer with $\text{TF}_{\text{pre}}$ (a one-layer two-head transformer), $\text{TF}_{\text{loop}}$ (a one-layer two-head transformer), and $\text{TF}_{\text{post}}$ (a two-layer two-head transformer), that outputs $\bm{w}^{i\top}_{T}\bm{x}_{i}$ for each input data $\bm{x}_{i}$, where $\bm{w}_{T}^{i}$ comes from $\arg\min_{\bm{w}}\mathcal{L}\left(\bm{w};\bm{P}^{i}\right)$ after $T$ steps of gradient descent. The emulation of each step is not exact, as there is some error introduced in each step. However, the error can be made arbitrarily close to zero by increasing the temperature of the softmax and adjusting another free parameter, neither of which affects the size of the network.
@@ 5 Experiments
In this section, we conduct a comprehensive empirical evaluation of the performance of AlgoFormer in tackling challenging tasks, specifically addressing regression with representation, AR(q) with representation, and CoT with MLPs, as outlined in Section 2. Additionally, we also implement AlgoFormer on the neural machine translation of German and English and AG News classification, demonstrating its expressiveness and effectiveness in real-world language tasks.
5.1 Experimental Settings and Hyperparameters
@@
Experimental settings. In all experiments, we adopt the decoder-based AlgoFormer, standard transformer (GPT-2), and vanilla looped transformer Yang et al. (2024). For synthetic tasks, we utilize $N=40$ in-context samples as input prompts and $d=20$ dimensional vectors with $D=256$ dimensional positional embeddings for all experiments. To ensure fairness in comparisons, all models are trained using the Adam optimizer, with a learning rate $\eta=1e-4$ and a total of 500K iterations to ensure convergence. The standard transformer is designed to have $L=12$ layers while pre-, looped and post-transformers are all implemented in one-layer. The default setting for the AlgoFormer, as well as the vanilla looped transformer, involves setting $(T,\Delta T)=(20,15)$. Here, the prompt formulation and the training loss in Equation (3) may slightly differ for different tasks. For example, in the AR(q) task, the input prompt is reformulated as $\bm{P}^{i}=[\bm{x}_{1},\cdots,\bm{x}_{i-1},\bm{x}_{i}]$ and $\bm{x}_{i+1}=f([\bm{x}_{i+1-q},\cdots,\bm{x}_{i}])$ is the target for prediction. However, the training strategy can be easily transmitted to other tasks. Here, both the iteration numbers $T_{0}$ and $T$ are hyperparameters, which will be analyzed in the next subsection.
Regression with representation. In this task, we instantiate a $3$-layer leaky ReLU MLPs, denoted as $\Phi^{*}(\cdot)$, which remains fixed across all tasks. The data generation process involves sampling a weight matrix $\bm{A}\in\mathbb{R}^{1\times 20}$. Subsequently, input-label pairs $\{(\bm{x}_{i},y_{i})\}_{i=1}^{40}$ are generated, where $\bm{x}_{i}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$, $\epsilon_{i}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ and $y_{i}=\bm{A}\Phi^{*}(\bm{x}_{i})+\epsilon_{i}$. In Figure 2a, we specifically set $\sigma=0$. Additionally, we explore the impact of different noise levels by considering $\sigma=0.1$ and $\sigma=1$. Here, our target is to learn the representation function $\Phi^{*}(\cdot)$ during training and to perform in-context learning of the weight matrix $\bm{A}$.
AR(q) with representation. For this task, we set $q=3$ and employ a $3$-layer leaky ReLU MLP denoted as $\Phi^{*}(\cdot)$, consistent across all instances. The representation function accepts a $60$-dimensional vector as input and produces $20$-dimensional feature vectors. The time series sequence $\{\bm{x}_{t}\}_{t=1}^{N}$ is generated by initially sampling $\bm{A}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20\times 20}\right)$. Then the sequence is auto-regressively determined, with $\bm{x}_{t+1}=\bm{A}\Phi^{*}\left([\bm{x}_{t+1-q},\cdots,\bm{x}_{t}]\right)+\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$. Here, our target is to learn the representation function $\Phi^{*}(\cdot)$ during training and to perform in-context learning of the weight matrix $\bm{A}$.
CoT with MLPs. In this example, we generate a 6-layer leaky ReLU MLP to serve as a CoT sequence generator, determining the length of CoT steps for each sample to be six. The CoT sequence, denoted as $\{\bm{x},\bm{s}^{1},\cdots,\bm{s}^{L}\}$ is generated by first sampling $\bm{x}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$, where $\bm{s}^{\ell}\in\mathbb{R}^{20}$ represents the intermediate state output from the $\ell$-th layer of the MLP. Here, our target is to perform in-context learning of the generator function (i.e., the 6-layer leaky ReLU MLP).
5.2 AlgoFormer Exhibits High Expressiveness